### v1

# Ahmed Heakl♠♣111 Equal Contributions , Abdullah Sohail111 Equal Contributions , Mukul Ranjan111 Equal Contributions , Rania Hossam111 Equal Contributions , Ghazi Ahmed Mohamed El-Geish , Omar Maher , Zhiqiang Shen , Fahad Khan♠♡ , Salman Khan♠♢

Abstract
With the growing adoption of Retrieval-Augmented Generation (RAG) in document processing, robust text recognition has become increasingly critical for knowledge extraction. While OCR (Optical Character Recognition) for English and other languages benefits from large datasets and well-established benchmarks, Arabic OCR faces unique challenges due to its cursive script, right-to-left text flow, and complex typographic and calligraphic features. We present KITAB-Bench, a comprehensive Arabic OCR benchmark that fills the gaps in current evaluation systems. Our benchmark comprises 8,809 samples across 9 major domains and 36 sub-domains, encompassing diverse document types including handwritten text, structured tables, and specialized coverage of 21 chart types for business intelligence. Our findings show that modern vision-language models (such as GPT-4, Gemini, and Qwen) outperform traditional OCR approaches (like EasyOCR, PaddleOCR, and Surya) by an average of $60\%$ in Character Error Rate (CER). Furthermore, we highlight significant limitations of current Arabic OCR models, particularly in PDF-to-Markdown conversion, where the best model Gemini-2.0-Flash achieves only $65\%$ accuracy. This underscores the challenges in accurately recognizing Arabic text, including issues with complex fonts, numeral recognition errors, word elongation, and table structure detection. This work establishes a rigorous evaluation framework that can drive improvements in Arabic document analysis methods and bridge the performance gap with English OCR technologies.
[Uncaptioned image]   KITAB-Bench: A Comprehensive Multi-Domain Benchmark for Arabic OCR and Document Understanding
Ahmed Heakl♠♣111 Equal Contributions , Abdullah Sohail111 Equal Contributions , Mukul Ranjan111 Equal Contributions , Rania Hossam111 Equal Contributions , Ghazi Ahmed Mohamed El-Geish , Omar Maher , Zhiqiang Shen , Fahad Khan♠♡ , Salman Khan♠♢ MBZUAI     Monta AI  Linköping University     Australian National University {ahmed.heakl,mabdullah.sohail,mukul.ranjan,salman.khan}@mbzuai.ac.ae {geish,omar}@monta.ai https://mbzuai-oryx.github.io/KITAB-Bench/
1 Introduction
Figure 1: Overview of the core domains and sub-domains in KITAB-Bench. Our benchmark spans nine major domains (e.g., OCR, charts to JSON, table recognition) and 36 sub-domains (e.g., scanned text, handwritten text, various chart types), providing a comprehensive evaluation framework for modern Arabic document processing and analysis.
With the upsurge in adoption of Retrieval-Augmented Generation (RAG) based systems for document processing, the quality of document ingestion pipelines has become increasingly critical. Optical Character Recognition (OCR) plays a crucial role in this pipeline, enabling the conversion of physical documents into machine-readable text and databases for enabling effective knowledge retrieval. Although significant progress has been made in the multilingual OCR JaidedAI (2020); Fu et al. (2024); Wei et al. (2024); Smith (2007), with comprehensive datasets like PubLayNet Zhong et al. (2019b), DocBank Li et al. (2020), M6Doc Cheng et al. (2023), and DocLayNet Pfitzmann et al. (2022), Arabic OCR continues to lag behind. This gap is largely due to the unique challenges of the Arabic script, including its cursive nature, complex typography, and right-to-left text orientation.
Figure 2: Overview of different tasks in our benchmark: Eight key components illustrating the task inputs and outputs for table recognition, chart understanding, text recognition, diagram analysis, VQA, line detection, layout analysis, and PDF-to-Markdown conversion, complete with input/output examples for each task.
Domain/ EXAMS-V Camel- MIDAD KHATT KITAB-
Characteristics Bench Bench (Ours)
PDF to Markdown
Layout Detection
Line Detection
Line Recognition
Table Recognition
Image to Text
Charts to JSON
Diagram to Code
VQA
Handwritten Samples
Open Source
Total Samples (#) 823 3,004 29,435 5,000 8,809
Table 1: Comparison of Arabic OCR Benchmarks Across Different Domains. Benchmarks compared: LaraBench Abdelali et al. (2023), CamelBench Ghaboura et al. (2024), MIDAD Bhatia et al. (2024), KHATT Mahmoud et al. (2014), and KITAB-Bench (Ours). (*: Only the Arabic samples are considered.) ($\dagger$: The test set of the dataset is considered.)
Existing Arabic OCR datasets (Table 1), like KHATT Mahmoud et al. (2014) and IFN/ENIT Pechwitz et al. (2002) focus mainly on handwritten text, whereas APTI Slimane et al. (2009) covers only specific aspects of printed text. These efforts fail to address advanced document processing challenges such as table parsing, font detection, and numeral recognition. Arabic benchmarks like CAMEL-Bench Ghaboura et al. (2024) and LAraBench Abdelali et al. (2023) evaluate large multimodal and language models, but they give limited attention to document understanding tasks. Consequently, there remains a need for a more comprehensive framework to systematically evaluate and compare Arabic OCR solutions. Our benchmark addresses these gaps by offering diverse document types and evaluation tasks to facilitate in-depth assessments of modern OCR systems.
We present KITAB-Bench, a comprehensive Arabic OCR benchmark spanning 9 domains and 36 sub-domains. Our framework evaluates layout detection (text blocks, tables, figures), multi-format recognition (printed/handwritten text, charts, diagrams), and structured output generation (HTML tables, DataFrame charts, markdown). This enables rigorous assessment of both basic OCR capabilities and advanced document understanding tasks.
The contributions of this work include (1) A comprehensive Arabic OCR benchmark covering multiple document types and recognition tasks. (2) Detailed evaluation metrics for assessing performance across different document understanding challenges. We also propose CharTeX and CODM metric to evaluate chart extraction and diagram extraction respectively. (3) Baseline results for popular OCR systems and Vision Language Models (VLMs), highlighting current limitations and areas for improvement. (4) A standardized framework for comparing Arabic OCR systems, facilitating future research and development.
Figure 3: Comparison of model performance across four document understanding tasks (Table Recognition, Image to Text, Diagram to JSON, and Layout Detection) showing successful and failed cases for different models including Ground Truth, EasyOCR, GPT-4, Qwen, Surya, Tesseract, Yolo, and DETR on Arabic document benchmark data.
2 Related Work
The development of robust Optical Character Recognition (OCR) systems has been extensively studied across document layout analysis Zhao et al. (2024); Shen et al. (2021); Paruchuri (2024b); JaidedAI (2020); Auer et al. (2024); Li et al. (2020), table detection Li et al. (2019); Paliwal et al. (2019); Nassar et al. (2022); Li et al. (2021); Schreiber et al. (2017), and document understanding Staar et al. (2018); Weber et al. (2023); Livathinos et al. (2021). While English OCR benefits from rich datasets like PubLayNet Zhong et al. (2019b), DocBank Li et al. (2020), M6Doc Cheng et al. (2023), and DocLayNet Pfitzmann et al. (2022), Arabic lacks standardized benchmarks for diverse fonts and layouts. Recent efforts like MIDAD Bhatia et al. (2024) curates extensive training data for Arabic OCR and handwriting recognition, while Peacock Alwajih et al. (2024) introduces culturally-aware Arabic multimodal models. Existing resources such as CAMEL-Bench Ghaboura et al. (2024), LAraBench Abdelali et al. (2023), MADAR Bouamor et al. (2018), OSACT Mubarak et al. (2022), and Tashkeela Zerrouki and Balla (2017) focus on language modeling or specific tasks rather than full-page OCR evaluation. Handwriting datasets including HistoryAr Pantke et al. (2014), IFN/ENIT Pechwitz et al. (2002), KHATT Mahmoud et al. (2014), APTI Slimane et al. (2009), and Muharaf Saeed et al. (2024) emphasize word/line recognition over document structure analysis.
Domain Total Samples
PDF to Markdown 33
Layout 2,100
Line Detection 378
Line Recognition 378
Table Recognition 456
Image to Text 3,760
Charts to DataFrame 576
Diagram to Json 226
VQA 902
Total 8,809
Table 2: Distribution of samples across different domains in our dataset. A more detailed count for different sub-domains and data sources is in Appendix A.
Arabic table recognition faces challenges from merged cells and RTL formatting Pantke et al. (2014). While methods like GTE Zheng et al. (2021), GFTE Li et al. (2021), CascadeTabNet Prasad et al. (2020), TableNet Paliwal et al. (2019), and TableFormer Nassar et al. (2022) advance Latin table detection, their effectiveness on Arabic documents remains unproven. Document conversion pipelines (CCS Staar et al. (2018), Tesseract Smith (2007), Docling Auer et al. (2024), Surya Paruchuri (2024b), Marker Paruchuri (2024a), MinerU Wang et al. (2024a), PaddleOCR Du et al. (2020)) lack Arabic-specific optimizations for segmentation and diacritic handling Mahmoud et al. (2018); Kiessling et al. (2019). This highlights the critical need for comprehensive Arabic OCR benchmarks addressing text recognition, table detection, and layout parsing.
3 KITAB-Bench
Our methodology offers a novel approach to benchmarking Arabic OCR systems via a comprehensive data collection strategy and a systematic evaluation framework. We gather curated samples from existing Arabic document datasets, manually collected and annotated PDFs, and employ a five-phase LLM-assisted human-in-the-loop pipeline (Figure 4) to generate diverse supplementary content. Our evaluation framework spans nine specialized tasks, enabling thorough assessment of OCR performance across various document processing challenges and providing a robust benchmark for Arabic document understanding tasks.
Figure 4: Synthetic Data Generation Pipeline: A 5-stage process using LLMs to generate topics, create raw data, produce visualization code, render charts, and perform human evaluation for quality control.
3.1 PDF Data Collection
We curated 33 diverse PDFs from online sources in academia, medicine, law, and literature. To ensure challenging cases, we selected documents featuring richly formatted tables with extensive color usage, merged cells, Arabic numerals, historical texts, watermarks, and handwritten annotations. Each PDF averaged three pages, and we then manually annotated them. This dataset comprehensively captures real-world complexities, making it a valuable benchmark for PDF-to-Markdown conversion.
3.2 LLM-Assisted Data Generation Pipeline
To generate data for charts, diagrams and tables, we implemented a five-phase LLM-assisted generation pipeline with human validation at critical stages, as illustrated in Figure 4. In Phase I (Topic Generation), our system employs an LLM to generate diverse topic names across multiple domains. This phase incorporates various personas (academic, legal, medical, technical) to ensure broad coverage of document types. Phase II (Data Generation) transforms the validated topics into structured raw data. The LLM generates content following Arabic linguistic and formatting conventions across various domains. In Phase III (Code Generation), the system converts the validated raw data into plotting code, with special attention to Arabic text rendering requirements and RTL content management. Phase IV (Image Rendering) utilizes specialized rendering engines (Mermaid, Plotly, Vegalite, HTML) to create visual representations while maintaining Arabic text integrity.
The final phase (Human Evaluation) implements rigorous quality control through expert validation. Evaluators filter charts, tables and diagrams based on detected anomalies and ensure adherence to Arabic-specific document conventions. This phase is crucial for maintaining the high quality of our benchmark dataset.
3.3 Dataset Statistics
Our benchmark dataset comprises over 8,809 samples across 9 major domains and 36 sub-domains, representing a comprehensive collection of Arabic document types for OCR evaluation. As detailed in Table 8, the dataset combines carefully curated samples from established datasets, manually annotation PDFs, and synthetically generated content created through our LLM-assisted pipeline (Figure 4). The Image-to-Text portion (3,760 samples) includes data from historical documents (HistoryAr Pantke et al. (2014)), handwritten text collections (Khatt Mahmoud et al. (2014), ADAB Boubaker et al. (2021), Muharaf Saeed et al. (2024)), and scene text (EvAREST Hassan et al. (2021)), while layout detection comprises 2,100 samples from BCE-Arabic-v1 Saad et al. (2016) and DocLayNet Pfitzmann et al. (2022).
For layout analysis, we incorporated 1,700 samples from BCE-Arabic-v1 dataset Saad et al. (2016), 400 samples from DocLayNet dataset Pfitzmann et al. (2022) focusing on financial, academic, legal, and patent documents. The line detection and recognition tasks contains 378 samples each from self-developed dataset. We further enriched the dataset with 500 samples from PATS-A01 El-Muhtaseb (2010) benchmark to ensure diverse representation.
Task Metric Surya Tesseract EasyOCR
Detection mAP@50 79.67 46.39 68.02
mAP@0.5:0.95 27.40 14.30 32.74
Recognition WER 1.01 1.00 0.53
CER 0.87 0.66 0.20
Table 3: Performance of different models on Line Detection and Line Recognition Task on our Benchmark
For handwritten text recognition, we assembled a comprehensive collection of 1,000 samples combining datasets from Khatt Mahmoud et al. (2014) (both paragraph and line-level annotations), Adab Boubaker et al. (2021), Muharaf Saeed et al. (2024), and OnlineKhatt Mahmoud et al. (2018). The benchmark also includes specialized content from ISI-PPT Wu and Natarajan (2017) (500 samples), and Hindawi Elfilali (2023) (200 samples) for various document types. Scene text understanding is supported by 800 samples from EvArest Hassan et al. (2021), providing real-world context diversity. A detailed table showing all the dataset is provided in the Appendix A.
A significant portion of our dataset consists of synthetically generated content, including 576 samples for Charts-to-DataFrame (spanning 16 different chart types), 422 samples for Diagram-to-Code (covering sequence diagrams, flowcharts, and tree maps), 456 samples for Tables-to-CSV/HTML, and 902 samples for VQA tasks. These synthetic samples were generated through our five-phase LLM-assisted human-in-the-loop pipeline (Figure 4). Every sample in our dataset - whether from existing sources or newly generated - underwent validation by native Arabic speakers before inclusion in the final benchmark. This rigorous validation, reinforced by expert review and automated checks, ensures high quality and authenticity across all domains. A detailed analysis is in Appendix C.
4 Experiments
Our experimental evaluation comprehensively assesses the capabilities of current OCR systems and state-of-the-art vision-language models (VLMs) across different Arabic and multilingual document understanding tasks. Figure 2 illustrates the nine distinct tasks in our evaluation framework.
We evaluate three categories of systems: VLMs, traditional OCR systems, and specialized document processing tools. For VLMs, we include both closed-source models like gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18 Hurst et al. (2024); Achiam et al. (2023), and gemini-2.0-flash Georgiev et al. (2024); Google DeepMind (2025), as well as open-source alternatives such as Qwen2-VL-7B Wang et al. (2024b), Qwen2.5-VL-7B Team (2025), and the AIN-7B Heakl et al. (2025). Traditional OCR approaches in our evaluation include Surya Paruchuri (2024b), Tesseract Smith (2007), EasyOCR JaidedAI (2020), and PaddleOCR Li et al. (2022); Du et al. (2021). For specialized document processing tasks, we employ systems like Docling Auer et al. (2024), and Marker Paruchuri (2024a). Layout detection capabilities are evaluated using methods implemented in Surya-layout Paruchuri (2024b), Yolo-doclaynet Zhao et al. (2024) from MinerU Wang et al. (2024a), and RT-DETR Zhao et al. (2023) based method in Docling Auer et al. (2024).
Dataset Metric Surya Yolo-doc- Detr
laynet (docling)
BCE mAP@0.5 0.506 0.470 0.750
mAP@0.5:0.95 0.381 0.369 0.566
Precision 0.751 0.608 0.626
Recall 0.593 0.592 0.725
F1 Score 0.635 0.585 0.654
DocLayNet mAP@0.5 0.675 0.404 0.758
mAP@0.5:0.95 0.469 0.335 0.541
Precision 0.782 0.527 0.635
Recall 0.856 0.503 0.770
F1 Score 0.799 0.499 0.670
Table 4: Performance comparison of layout detection models using different evaluation metrics
4.1 Evaluation Frameworks and Metrics
Our evaluation framework comprises nine specialized tasks designed to assess different aspects of Arabic OCR systems, as demonstrated in Figure 2. Each task addresses specific challenges in Arabic document processing. For this reason, we employ task-specific metrics to evaluate different aspects of document understanding.
PDF-to-Markdown: It evaluates the conversion of Arabic PDFs to structured markdown while preserving the text and table structure. Since both table and text structure are important, for evaluating PDF to Markdown conversion quality, we propose MARS (Markdown Recognition Score), which combines chrF Popović (2015) with Tree-Edit-Distance-based Similarity (TEDS) Zhong et al. (2020) :
$\text{MARS}=\alpha\cdot\text{chrF}_{3}+(1-\alpha)\cdot\text{TEDS}(T_{a},T_{b})$ (1)
where $\alpha$ ($0\leq\alpha\leq 1$) is the weight. $T_{a}$ represent predicted table structure and $T_{b}$ the ground truth structure.
Table Recognition: We evaluate table extraction using both HTML and CSV formats, where HTML format (evaluated using TEDS Zhong et al. (2020)) preserves rich structural information including cell spans and hierarchical relationships crucial for complex Arabic tables, while CSV format (evaluated using Jaccard Index 2) focuses on raw data extraction optimized for machine processing and data analysis pipelines. This dual-format evaluation ensures systems can both maintain complex table structures for human readability and provide clean, structured data for automated processing, specifically important for RAG based systems.
$J(P,G)=\frac{|P\cap G|}{|P\cup G|}=\frac{|P\cap G|}{|P|+|G|-|P\cap G|}$ (2)
where $|P\cap G|$ represents the number of exact matching cells between predicted and ground truth tables, and $|P\cup G|$ represents the total number of unique cells across both tables.
Chart-to-Dataframe: This task evaluates extracting structured data from Arabic charts into machine-readable dataframes. Systems must accurately parse numerical values, text labels, and preserve data relationships across chart types (bar, line, pie). We use the Structuring Chart-oriented Representation Metric (SCRM) Xia et al. (2024)—which combines type recognition, topic understanding, and structural numerical fidelity (see Appendix D)—and also propose our own CharTeX (Chart Extraction Score) metric. CharTeX combines the chrF scores for chart type and topic with the jaccord index for the dataframe, using fuzzy matching (80% threshold) when columns do not exactly align.
$\text{Metric}=\alpha J_{type}+\beta J_{topic}+(1-\alpha-\beta)J_{data}$ (3)
Here, $J_{type}$ and $J_{topic}$ denote the chrF scores between the predicted and ground-truth chart type and topic, while $J_{data}$ measures the structural similarity of the predicted and ground-truth JSON data.
Diagram-to-JSON: This task evaluates the conversion of Arabic flowcharts and technical diagrams into JSON while preserving semantic relationships and technical specifications. We propose CODM (Code-Oriented Diagram Metric), extending SCRM Xia et al. (2024), with the same fomulation as in Eq 3. More detail about this metric is provided in Appendix E.
Table Extraction End-to-End PDF
Model Group Models TEDS (HTML) Jaccard (CSV) CHrF (Text) TEDS (Table) MARS
Closed GPT-4o 85.76 66.36 69.62 60.61 65.12
GPT-4o-mini 69.32 49.50 56.59 52.69 54.64
Gemini-2.0-Flash 83.08 65.55 75.75 55.55 65.65
Open Qwen2-VL-7B 57.83 40.20 40.30 2.54 21.42
Qwen2.5-VL-7B 59.31 59.58 69.21 11.65 40.43
AIN-7B 75.94 64.83 56.52 49.32 52.92
Framework Tesseract 28.23D 38.64I 14.85D 16.04I 59.91D 45.44D 52.68D
EasyOCR 49.10D 39.09I 23.83D 17.88I 57.46D 51.12D 54.29D
Surya 50.15M 70.42M 58.38M 44.29M 51.34M
DDocling Auer et al. (2024) pipeline  IImg2Table Cattan (2021) pipeline  MMarker Paruchuri (2024a) pipeline
Table 5: Performance comparison of different models for table extraction and end-to-end PDF to markdown conversion tasks on our benchmark.
Image-to-Text: This task assess the basic text recognition capabilities across different Arabic fonts and styles, including the handling of cursive script connections, diacritical marks, and various text orientations. We use we use Character Error Rate (CER) and Word Error Rate (WER). For a predicted text sequence $\hat{y}$ and ground truth sequence $y$, CER is computed as: $\text{CER}=\frac{\text{L}(y,\hat{y})}{|y|}$, where $\text{L}(y,\hat{y})$ is the Levenshtein distance between character sequences and $|y|$ is the ground truth length. WER is calculated the same way with words as the unit of error.
Visual Question Answering: Tests the ability of models to understand and reason about Arabic document content, we evaluate using standard accuracy for MCQ questions and exact word match.
Line Detection: Focuses on the accurate identification and processing of individual text lines in Arabic documents. We evaluate using mean Average Precision (mAP) at different Intersection over Union (IoU) thresholds: mAP@0.5 and mAP@0.5:0.95, which assess the localization accuracy of detected text lines.
Layout Detection: Assesses document structure analysis capabilities, including the identification of headers, paragraphs, and complex layout elements in Arabic documents. Performance is measured using mAP@0.5 and mAP@0.5:0.95 for localization accuracy, complemented by Precision, Recall, and F1 scores to evaluate the overall detection quality across different layout components.
All metrics are computed on our diverse benchmark dataset, which encompasses various document types and complexity levels in both Arabic and multilingual contexts. Table 10 provides a detailed mapping of tasks, metrics, and evaluated systems.
Group Models CHrF $\uparrow$ CER $\downarrow$ WER $\downarrow$
Closed GPT-4o 61.01 0.31 0.55
GPT-4o-mini 47.21 0.43 0.71
Gemini-2.0-Flash 77.95 0.13 0.32
Open Qwen2VL-7B 33.94 1.48 1.55
Qwen2.5VL-7B 49.23 1.20 1.41
AIN-7B 78.33 0.20 0.28
Framework Tesseract 39.62 0.54 0.84
EasyOCR 45.47 0.58 0.89
Paddle 16.73 0.79 1.02
Surya 20.61 4.95 5.61
Table 6: Performance comparison of models for OCR (image to text) tasks on our benchmark. A detailed performance comparison among different open-source dataset is available in Appendix B
4.2 Experimental Setup
We implement our evaluation pipeline with careful consideration of hyperparameters for different metrics. All experiments use NVIDIA A100 GPUs. For VLMs, we use their official implementations or API endpoints. Traditional OCR systems are evaluated using pre-trained models provided by the frameworks. For PDF-to-Markdown evaluation metric MARS 1, we choose $\alpha=0.5$ and $\alpha=0.5$ and $\beta=0.2$ for Diagram-to-JSON evaluation metric CODM. We average the results over multiple runs, with performance comparisons shown in different tables [3, 6, 5, 7, and 4].
Group Model Chart Diagram Visual QA
SCRM CharTeX CODM MTVQAO ChartsVQAM DiagramsVQAM PATDVQAM Average
Closed GPT-4o 68.6 45.95 61.6 32.00 77.00 85.29 82.50 69.19
GPT-4o-mini 67.2 43.33 61.4 26.80 58.00 83.33 80.00 62.03
Gemini-2.0-Flash 71.4 56.28 71.8 35.00 72.00 88.24 75.50 67.68
Open Qwen2-VL-7B 56.6 21.59 63.0 19.60 59.00 82.35 77.50 59.61
Qwen2.5-VL-7B 36.2 22.08 59.2 23.00 74.00 79.41 74.50 62.72
AIN-7B 66.6 34.61 66.40 31.50 75.00 85.29 87.00 69.69
Table 7: Model Performance on Chart Understanding, Diagram Parsing, and Visual Question Answering Tasks. For VQA tasks, $O$ denotes open-ended question type from MTVQA Tang et al. (2024) dataset and $M$ denotes MCQ type questions.
5 Results and Discussion
In this section, we present a comprehensive evaluation of different models across different tasks of our framework. The results provide a clear distinction between the performance of closed-source models, open-source models, and framework-based solutions, revealing both their strengths and limitations. We observe very clear performance gap between closed and open-source solutions. While closed-source models like Gemini-2.0-Flash consistently outperform other models almost all the tasks.
5.1 Charts, Diagrams, and VQA
Table [7] presents model performance across different chart and diagram understanding tasks, evaluated using SCRM and CharTeX (for charts), and VQA-based accuracy metrics. Among closed-source models, Gemini-2.0 achieves the highest performance on chart understanding metrics, scoring 71.4% on SCRM and 56.28% on CharTeX. The performance gap between Gemini-2.0 and GPT-4o is particularly pronounced in CharTeX evaluation (10.33%) compared to SCRM (2.8%). Open-source models shows a significant limitation in complex chart understanding. While their SCRM scores remain competitive, both Qwen variants score below 23% on CharTeX evaluation. The visual question-answering results reveal an important exception to the general closed-source advantage. AIN achieves 87% on PATDVQA, surpassing Gemini-2.0 by 11.5%. AIN also shows competitive performance on MTVQA (31.50%), which is similar to GPT-4o and  4% better than GPT-4o-mini. This shows that open-source models can be competitive with closed-source alternatives.
5.2 Layout and Lines: Document Structure
Our evaluation of document structure understanding reveals distinct performance patterns across layout detection and line processing tasks. In layout detection (Table 4), RT-DETR Zhao et al. (2023) achieves superior overall performance with mAP@0.5 scores of 0.750 and 0.758 on BCE (arabic only) and DocLayNet (english) datset respectively. However, Surya Paruchuri (2024b) demonstrates higher precision (0.782 on DocLayNet, 0.751 on BCE), despite lower recall rates. This trade-off suggests that different architectures optimize for different aspects of layout detection.
The line processing results (Table 3) highlight a clear contrast between detection and recognition capabilities. While Surya excels in detection with a mAP@0.50 of 79.67%, EasyOCR demonstrates superior recognition performance (WER: 0.53, CER: 0.20). This inverse relationship between detection and recognition performance across models indicates a fundamental challenge in optimizing both capabilities simultaneously. Notably, Tesseract shows consistent but lower performance across both metrics, suggesting that newer architectures have made significant improvements over traditional approaches. We also observe that no single model excels at both detection and recognition, which requires for hybrid solutions.
5.3 Tables, OCR, and PDF-to-Markdown
Across table extraction tasks (Table 5), closed-source models maintain a clear advantage, with GPT-4o achieving 85.76% TEDS and 66.36% Jaccard scores. Among open-source models, AIN (75.94% TEDS) significantly outperforms Qwen variants, while specialized frameworks like Surya achieve competitive results (70.42% Jaccard) through targeted pipelines. In OCR evaluation (Table 6), Gemini-2.0-Flash leads with the lowest error rates (CER: 0.13, WER: 0.32). Notably, AIN matches this performance level (WER: 0.28), while traditional OCR frameworks like EasyOCR and Tesseract show moderate performance (CER: 0.58, 0.54). The significant performance drop in Paddle (CER: 0.79) and Surya (CER: 4.95) highlights the challenges in developing robust OCR systems.
End-to-end document processing (Table 5) reveals the largest gaps between approaches. Closed-source models maintain consistent performance (GPT-4o: 65.12% MARS, Gemini-2.0: 65.65% MARS), while open-source models show substantial degradation (Qwen2-VL-7B: 21.42% MARS). Framework approaches achieve better stability, with Tesseract and EasyOCR scoring above 50% MARS, suggesting that specialized pipelines can partially bridge the gap with larger models in complete document processing tasks.
Our comprehensive evaluation demonstrates that while closed-source models maintain superior performance over open-source models across most tasks, specialized frameworks like Surya, RT-DETR Layout, and EasyOCR achieve competitive performance in targeted scenarios like table extraction, layout detection, and text recognition respectively. However, this framework advantage significantly diminishes in end-to-end pdf-to-markdown tasks where the integration capabilities of large models prove crucial, as evidenced by the performance gaps between commercial VLMs and traditional systems like EasyOCR, Surya and Tesseract in End-to-End PDF task (Table 5).
6 Conclusion
We introduce a comprehensive benchmark for Arabic OCR that fills the gap in standardized evaluation frameworks for Arabic document processing. Our dataset of 8,809 samples across nine major domains is the most diverse collection assembled for OCR evaluation, incorporating handwritten, scanned, synthetic, and scene text, as well as complex tables, charts, and end-to-end pdf-to-markdown. This framework extends beyond simple text recognition to include structural document analysis and enables systematic assessment of OCR performance across various fonts, styles, and layouts.
7 Limitations and Future Directions
Despite its contributions, this benchmark has limitations. While it covers diverse Arabic document types, it lacks full representation of historical manuscripts and low-resource dialects. Future work should expand to include these, along with scanned records from government, academic, and financial institutions.
Another key limitation is in table and chart recognition, where OCR models struggle with structure preservation, header detection, and merged cell parsing. Though our benchmark introduces challenges in these areas, further refinements are needed for robust multimodal OCR capable of jointly processing text, tables, and figures. Future advancements should focus on dataset expansion, novel evaluation metrics, deep learning refinements, and cross-lingual OCR innovations to enhance Arabic VLMs.
Appendix A Source of the Existing Dataset Collection
Our benchmark integrates diverse data sources to ensure comprehensive coverage of Arabic document types. As detailed in Table 2, the dataset combines manually curated samples, synthetic data generated through our LLM-assisted pipeline (Figure 4), and existing publicly available datasets. Key sources include:
  • Handwritten Text: KHATT (paragraph and line-level annotations), ADAB, Muharaf, and OnlineKhatt.
  • Historical Documents: HistoryAr and HistoricalBooks.
  • Scene Text: EvAREST for real-world context diversity.
  • Layout Analysis: BCE-Arabic-v1 and DocLayNet.
  • Synthetic Content: 576 chart samples (16 types) and 422 diagram samples generated via our five-phase pipeline (Section 3.2).
The dataset emphasizes domain diversity, covering academic, medical, legal, financial, and technical documents. All samples underwent rigorous validation by native Arabic speakers to ensure linguistic and structural accuracy.
Appendix B Detailed Performance Comparison
Table 9 provides granular performance metrics for VLMs and OCR frameworks across 12 Arabic text recognition datasets. Gemini-2.0-Flash demonstrates exceptional robustness on synthetic datasets (CER: 0.01 on PATS), while AIN-7B excels in historical manuscript recognition (CER: 0.26 on HistoryAr). Traditional OCR systems like Tesseract show limitations in handwritten text (CER: 1.26 on HistoryAr), highlighting the need for script-specific optimizations.
Appendix C Data Analysis
Our data generation pipeline (Figure 4) enabled the creation of 1,502 synthetic samples (576 charts, 422 diagrams, 456 tables). The pipeline’s human validation phase rejected 18% of initial outputs due to RTL formatting errors or semantic inconsistencies. As shown in Figure 5 and 6, domain-specific prompts ensured adherence to Arabic linguistic conventions during LLM-assisted generation. The final dataset exhibits balanced representation across:
  • Font Styles: 21 Arabic calligraphic styles
  • Document Types: 36 sub-domains including financial reports and technical manuals
  • Structural Complexity: 43% of tables contain merged cells; 29% of charts use dual-axis configurations
Appendix D Tasks Models and Metrics
Table 10 maps evaluation tasks to corresponding models and metrics. The framework evaluates nine core capabilities:
  • Structural Understanding: Layout detection (mAP), line detection (IoU)
  • Content Extraction: Text recognition (CER), table parsing (TEDS)
  • Semantic Reasoning: VQA accuracy, chart-to-dataframe conversion (SCRM)
  • Specialized metrics like MARS ( $\alpha$=0.5) address the dual requirements of text fidelity and structural preservation in PDF-to-Markdown conversion.
Appendix E SCRM and CODM
The Structuring Chart-oriented Representation Metric (SCRM) evaluates chart understanding through three components:
$\text{SCRM}=0.4J_{\text{type}}+0.3J_{\text{topic}}+0.3J_{\text{data}}$ (4)
where $J_{type}$, and $J_{topic}$ are chrF scores, and $J_{data}$ measures JSON structural similarity.
The Code-Oriented Diagram Metric (CODM) extends SCRM for flowcharts and technical diagrams:
$\text{CODM}=0.5J_{\text{topology}}+0.5J_{\text{semantics}}$ (5)
assessing both node-edge relationships and semantic labels. As shown in Figure 5 and 6, domain-specific prompts guided model responses for metric calculation. For instance, sequence diagrams required strict adherence to Arabic UML notation standards during evaluation.
Domain Sub-Domain Dataset Source Original Selected Total
PDF to Markdown General Manual 33 33 33
Layout Detection Docs BCE-Arabic-v1 Saad et al. (2016) 1.9k 1,700 2,100
DocLayNet Pfitzmann et al. (2022) 80k 400
Line Detection Docs Manual 375 378 378
Line Recognition Docs Manual 375 378 378
Table Recognition Financial Pixmo Deitke et al. (2024) 490 456 456
Image to Text Synthetic PATS El-Muhtaseb (2010) 21.6k 500 3,760
SythenAR 39.1k 500
Historical HistoryAr Pantke et al. (2014) 1.5k 200
HistoricalBooks 40 10
Hand. Paragraph Khatt Mahmoud et al. (2014) 2.72k 200
Hand. Word ADAB Boubaker et al. (2021) 15k 200
Hand. Line Muharaf Saeed et al. (2024) 24.5k 200
OnlineKhatt Mahmoud et al. (2018) 8.5k 200
Khatt Mahmoud et al. (2014) 13.4k 200
PPT ISI-PPT Wu and Natarajan (2017) 86.5k 500
Blogs ArabicOCR 20.3k 50
Hindawi Elfilali (2023) 79k 200
Scene EvAREST Hassan et al. (2021) 5.59k 800
Charts to DataFrame Bar Synthetic 100 61 576
Line Synthetic 100 43
Pie Synthetic 100 56
Box Synthetic 100 31
Violin Synthetic 100 36
Area Synthetic 50 29
SunBurst Synthetic 30 15
Dot Synthetic 30 15
Dual Axis Synthetic 20 26
Density Curve Synthetic 10 5
Bubble Synthetic 20 13
Grouped Bar Synthetic 50 60
Stacked Bar Synthetic 50 82
Histogram Synthetic 100 70
HeatMap Synthetic 10 11
Scatter Synthetic 100 23
Diagram to Json Sequence Synthetic 50 46 226
Funnel Synthetic 20 52
Class Synthetic 20 30
Network Synthetic 20 18
Venn Synthetic 20 7
FlowChart Synthetic 100 112
TreeMap Synthetic 100 157
VQA Diagrams Manual 102 102 902
Charts Manual 105 100
News Letter PATD Bouressace and Csirik (2019) 2.42k 200
Scene MTVQA 818 500
Total Dataset Size 8,809
Table 8: Dataset Distribution Across Different Domains, sub-domains and Data Source
Dataset Size GPT-4o GPT-4o-mini Gemini-2.0-Flash Qwen2-VL
CER WER CER WER CER WER CER WER
PATS 500 0.23 0.30 0.53 0.71 0.01 0.02 1.02 1.02
SythenAR 500 0.09 0.20 0.14 0.32 0.07 0.17 0.59 1.13
HistoryAr 200 0.51 0.82 0.67 0.96 0.28 0.64 3.46 2.86
HistoricalBooks 10 0.41 0.76 0.59 0.88 0.05 0.22 1.90 2.16
Khatt 200 0.45 0.74 0.64 0.91 0.19 0.45 1.12 5.04
Adab 200 0.30 0.73 0.35 0.83 0.19 0.56 0.63 1.08
Muharaf 200 0.56 0.90 0.63 0.94 0.33 0.69 3.57 2.87
OnlineKhatt 200 0.29 0.63 0.41 0.76 0.17 0.44 1.30 2.01
ISI-PPT 500 0.08 0.18 0.15 0.31 0.06 0.15 1.03 1.06
ArabicOCR 50 0.06 0.26 0.16 0.46 0.00 0.02 1.25 1.50
Hindawi 200 0.34 0.56 0.48 0.71 0.01 0.04 1.82 2.05
EvArest 800 0.20 0.38 0.25 0.51 0.18 0.36 0.41 0.95
3,760 0.31 0.55 0.43 0.71 0.13 0.32 1.48 1.20
Dataset Size Qwen2.5-VL AIN Tesseract Surya
CER WER CER WER CER WER CER WER
PATS 500 0.26 0.36 0.00 0.00 0.14 0.28 4.66 4.67
SythenAR 500 0.21 0.40 0.04 0.16 0.31 0.72 4.82 7.90
HistoryAr 200 0.47 0.83 0.26 0.54 0.72 1.26 10.32 12.78
HistoricalBooks 10 0.33 0.72 0.84 0.88 0.74 0.99 6.81 6.30
Khatt 200 0.07 0.22 0.61 1.12 0.67 1.06 4.25 3.77
Adab 200 0.00 0.01 1.00 1.00 1.00 1.14 7.28 8.71
Muharaf 200 0.61 0.96 0.38 0.54 0.77 1.22 6.19 7.48
OnlineKhatt 200 0.36 0.70 0.03 0.12 0.59 1.20 6.71 6.95
ISI-PPT 500 0.36 0.54 0.52 0.53 0.31 0.64 4.25 3.77
ArabicOCR 50 1.00 1.00 0.01 0.01 0.01 0.01 2.75 3.58
Hindawi 200 1.00 1.00 0.11 0.15 0.31 0.72 0.15 0.20
EvArest 800 0.19 0.36 0.30 0.32 0.85 1.02 5.91 3.86
3,760 0.28 0.54 0.20 0.58 0.89 0.79 4.95 5.61
Table 9: Performance comparison of Large Vision-Language Models on KITAB-Bench (lower is better).
Task Metrics Open LLMs Closed LLMs OCR Systems
Document Understanding Tasks
PDF to Markdown chrF + TEDS Docling Marker MinerU PDF-Extract-Kit
Layout Detection mAP@0.5 mAP@0.5:0.95 Precision Recall F1 Surya Yolo-doclaynet (MinerU) Detr (docling)
Line Detection mAP@0.5 mAP@0.5:0.95 Surya Tesseract EasyOCR
Line Recognition WER, CER Surya Tesseract EasyOCR
Table Understanding Tasks
Tables Recognition (HTML) TEDS Zhong et al. (2019a) Qwen2-VL Qwen2.5-VL AIN PaliGemma GPT-4o GPT-4o-mini Gemini-2.0-Flash Docling[EasyOCR] Docling[Tesseract] Marker Img2Table[EasyOCR] Img2Table[Tesseract]
Tables Recognition (CSV) Jaccard Index Qwen2-VL Qwen2.5-VL AIN PaliGemma GPT-4o GPT-4o-mini Gemini-2.0-Flash Docling[EasyOCR] Docling[Tesseract] Marker Img2Table[EasyOCR] Img2Table[Tesseract]
Visual Understanding Tasks
Image to Text CER, WER chrF, BLEU METEOR Qwen2-VL Qwen2.5-VL AIN-7B PaliGemma GPT-4o GPT-4o-mini Gemini-2.0-Flash Docling[EasyOCR] Docling[Tesseract] Marker Img2Table[EasyOCR] Img2Table[Tesseract]
Charts to DataFrame SCRM Xia et al. (2024, 2023) Qwen2-VL Qwen2.5-VL AIN PaliGemma GPT-4o GPT-4o-mini Gemini-2.0-Flash
Diagram to Json SCRM Qwen2-VL Qwen2.5-VL AIN-7B PaliGemma GPT-4o GPT-4o-mini Gemini-2.0-Flash
VQA Accuracy + Word Match Score Qwen2-VL Qwen2.5-VL AIN-7b PaliGemma GPT-4o GPT-4o-mini Gemini-2.0-Flash
Table 10: Comprehensive evaluation metrics and models for document understanding tasks. The table is organized into three main categories: document understanding, table understanding, and visual understanding tasks. Each task is evaluated using specific metrics and implemented across various models and OCR systems.
Figure 5: Prompts for Different Task Categories.
Figure 6: Prompts for Diagrams and Tables.